OpenStudy (anonymous):

y = -2x - 3 please help me. step by step how to solve this.. :(

OpenStudy (anonymous):

hELP

OpenStudy (anonymous):

look ! This is the form : y=mx+d Now at there m is behind x .! It s -2

OpenStudy (anonymous):

the answer is -2.. :/ still confused

OpenStudy (anonymous):

Yeah !

OpenStudy (uri):

What do you have to do? Find slope?

OpenStudy (uri):

y = -2x - 3 formula is y=mx+b -2 is the slope.